# Cross-correlation between the curvature perturbations and magnetic fields in pure ultra slow roll inflation

### v1

@@ IV.2 Three-point cross-correlation
We shall now turn to the computation of the three-point cross-correlation using the interaction Hamiltonian (48). In real space, the cross-correlation between the curvature perturbation and magnetic fields is defined as
@@
where $\hat{H}_{\mathrm{int}}$ is the operator associated with the interaction Hamiltonian (48), and $\eta_{\mathrm{i}}$ denotes the time when the initial conditions are imposed on the perturbations. Moreover, the square brackets denote the commutator between the operators denoting modes of the curvature perturbation, the magnetic field, and the Hamiltonian operator.
To describe the cross-correlation between the modes of the curvature perturbation and the magnetic field, we shall introduce the function ${\mathcal{B}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ that is defined through the relation [^2] In the definition (52) of the three-point function, we have introduced an additional factor of $(4\,\pi)$ in comparison to the earlier works in the literature (in this context, see Refs. [88, 87, 89, 90, 101]). Note that our electromagnetic actions (16) and (34)) contain an additional factor of $1/(4\,\pi)$ in comparison to the action in the previous works. In order to ensure that the expression for the power spectrum of the magnetic field remains the same, we had included an overall factor of $\sqrt{4\,\pi}$ in the expression (20) for the vector potential $A_{i}$. However, such a definition for $A_{i}$ introduces another factor of $(4\,\pi)$ when we evaluate the quantity on the right hand side of Eq. (51), which is not present if we instead follow the normalization conventions used in the aforementioned works. Therefore, in order to ensure that the quantity ${\mathcal{B}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ remains the same across the different conventions, we have introduced the factor of $(4\,\pi)$ in Eq. (52).
@@
On making use of the expression (51), the interaction Hamiltonian (48), the definition (52) and, finally, Wick’s theorem, we find that the function ${\mathcal{B}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ can be written as follows [88, 87, 89, 90, 99, 101]:
@@
${\mathcal{G}}_{2}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=i\,\left[k_{2}^{2}\,k_{3}^{2}+(\bm{k}_{2}\cdot\bm{k}_{3})^{2}\right]\,\int_{\eta_{\mathrm{i}}}^{\eta_{\mathrm{e}}}{\mathrm{d}}\eta\,J^{2}(\eta)\,f_{k_{1}}^{\ast}(\eta)\,\bar{A}_{k_{2}}^{\ast}(\eta)\,\bar{A}_{k_{3}}^{\ast}(\eta),$ (54b)
${\mathcal{G}}_{3}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$ $=2\,i\,(1+2\,n)\,(\bm{k}_{2}\cdot\bm{k}_{3})\,\int_{\eta_{\mathrm{i}}}^{\eta_{\mathrm{e}}}\frac{{\mathrm{d}}\eta}{a\,H}\,J^{2}(\eta)\,f_{k_{1}}^{\prime\ast}(\eta)\,\bar{A}_{k_{2}}^{\prime\ast}(\eta)\,\bar{A}_{k_{3}}^{\prime\ast}(\eta),$ (54c)
${\mathcal{G}}_{4}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-i\,(1-2\,n)\,\left[k_{2}^{2}\,k_{3}^{2}+(\bm{k}_{2}\cdot\bm{k}_{3})^{2}\right]\,\int_{\eta_{\mathrm{i}}}^{\eta_{\mathrm{e}}}\frac{{\mathrm{d}}\eta}{a\,H}\,J^{2}(\eta)\,f_{k_{1}}^{\prime\ast}(\eta)\,\bar{A}_{k_{2}}^{\ast}(\eta)\,\bar{A}_{k_{3}}^{\ast}(\eta),$ (54d)
${\mathcal{G}}_{5}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=i\,\left[(\bm{k}_{1}\cdot\bm{k}_{2})\,k_{3}^{2}+(\bm{k}_{1}\cdot\bm{k}_{3})\,(\bm{k}_{2}\cdot\bm{k}_{3})\right]\,\int_{\eta_{\mathrm{i}}}^{\eta_{\mathrm{e}}}\frac{{\mathrm{d}}\eta}{a\,H}\,J^{2}(\eta)\,f_{k_{1}}^{\ast}(\eta)\,\bar{A}_{k_{2}}^{\prime\ast}(\eta)\,\bar{A}_{k_{3}}^{\ast}(\eta)$
$\quad+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$ (54e)
${\mathcal{G}}_{6}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{i}{k_{1}^{2}}\,\left[(\bm{k}_{1}\cdot\bm{k}_{2})\,k_{3}^{2}+(\bm{k}_{1}\cdot\bm{k}_{3})\,(\bm{k}_{2}\cdot\bm{k}_{3})\right]\,\int_{\eta_{\mathrm{i}}}^{\eta_{\mathrm{e}}}\frac{{\mathrm{d}}\eta}{a}\,J^{2}(\eta)\,\epsilon_{1}(\eta)\,f_{k_{1}}^{\prime\ast}(\eta)\,\bar{A}_{k_{2}}^{\prime\ast}(\eta)\,\bar{A}_{k_{3}}^{\ast}(\eta)$
$\quad+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3}.$ (54f)
@@ IV.3 Non-Gaussianity parameter
$\hat{B}_{i\,{\bm{k}}}(\eta)$ $=\hat{B}_{i\,{\bm{k}}}^{{}_{\mathrm{G}}}(\eta)+\frac{1}{(2\,\pi)^{3/2}}\int{\mathrm{d}}^{3}\bm{p}\,\hat{{\mathcal{R}}}_{\bm{p}}^{{}_{\mathrm{G}}}(\eta)\,\hat{B}_{i\,{\bm{k}-\bm{p}}}^{\rm G}(\eta)$
@@ V.1 Power spectra
Clearly, the amplitude of $\mathcal{P}_{{}_{\mathrm{B}}}(k)$ in ultra slow roll inflation is suppressed by a factor of $(\eta_{\mathrm{e}}/\eta_{\mathrm{i}})^{12}\sim 10^{-300}$. Such a small value of $\mathcal{P}_{{}_{\mathrm{B}}}(k)$ will not be sufficient to serve as seeds for magnetic fields with the strengths observed in the inter-galactic medium today. Despite these shortcomings of the scenario of pure ultra slow roll, we pursue the analysis with the aim of eventually utilizing the coupling function that we have constructed in a realistic model of inflation with a brief epoch of ultra slow roll. We shall now proceed with these parameter values to calculate and examine the amplitude and shape of the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$.
@@ V.2.1 Case of $p=6$
${\mathcal{B}}_{2}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{9\,H_{{}_{\mathrm{I}}}^{6}}{64\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{6}\,\left[k_{1}^{4}+k_{2}^{4}+k_{3}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+6\,k_{2}^{2}\,k_{3}^{2}\right],$ (63b)
${\mathcal{B}}_{3}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{3\,H_{{}_{\mathrm{I}}}^{6}\,\eta_{\mathrm{i}}^{2}}{32\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{1}^{3}\,k_{2}^{3}\,k_{3}^{3}\,\epsilon_{1\mathrm{i}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{4}\,(k_{1}^{2}-k_{2}^{2}-k_{3}^{2})\,\left[k_{1}^{3}-4\,(k_{2}^{3}+k_{3}^{3})\right],$ (63c)
${\mathcal{B}}_{4}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{63\,H_{{}_{\mathrm{I}}}^{6}}{64\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{6}\,\left[k_{1}^{4}+k_{2}^{4}+k_{3}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+6\,k_{2}^{2}\,k_{3}^{2}\right],$ (63d)
${\mathcal{B}}_{5}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$ $=\frac{9\,H_{{}_{\mathrm{I}}}^{6}\,\eta_{\mathrm{i}}^{2}}{128\,M_{{}_{\mathrm{Pl}}}^{2}\,\epsilon_{1\mathrm{i}}\,k_{1}^{3}\,k_{2}^{3}\,k_{3}^{5}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{4}\,\left[2\,k_{3}^{2}\,(k_{1}^{2}+k_{2}^{2})+(k_{1}^{2}-k_{2}^{2})^{2}-3\,k_{3}^{4}\right]\,(3\,k_{1}^{3}+4\,k_{2}^{3})+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$
${\mathcal{B}}_{6}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-\frac{27\,H_{{}_{\mathrm{I}}}^{7}}{16\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{1}^{5}\,k_{2}^{3}\,k_{3}^{5}\,(k_{1}+k_{2}+k_{3})^{6}\,\eta_{\mathrm{e}}^{3}}\,\left[k_{1}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}-k_{3}^{2})+k_{2}^{4}+2\,k_{2}^{2}\,k_{3}^{2}-3\,k_{3}^{4}\right]$
×$\quad\times\biggl{[}4\,k_{1}^{5}+24\,k_{1}^{4}\,(k_{2}+k_{3})+8\,k_{1}^{3}\,(6\,k_{2}^{2}+15\,k_{2}\,k_{3}+7\,k_{3}^{2})$
$\quad+k_{1}^{2}\,(43\,k_{2}^{3}+168\,k_{2}^{2}\,k_{3}+216\,k_{2}\,k_{3}^{2}+56\,k_{3}^{3})+6\,k_{1}\,(3\,k_{2}^{4}+15\,k_{2}^{3}\,k_{3}+28\,k_{2}^{2}\,k_{3}^{2}+20\,k_{2}\,k_{3}^{3}+4\,k_{3}^{4})$
$\quad+(k_{2}+k_{3})^{2}\,(3\,k_{2}^{3}+12\,k_{2}^{2}\,k_{3}+16\,k_{2}\,k_{3}^{2}+4\,k_{3}^{3})\biggr{]}+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3}.$ (63e)
Clearly, among the different contributions, it is the second and the fourth terms, viz. ${\cal B}_{2}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ and ${\cal B}_{4}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$, that dominate. Note that these contributions behave as $(\eta_{\mathrm{i}}/\eta_{\mathrm{e}})^{6}$, just as the scalar power spectrum $\mathcal{P}_{{}_{\mathrm{S}}}(k)$ does when $p=6$ [cf. Eq. (13)]. Since the magnetic power spectrum $\mathcal{P}_{{}_{\mathrm{B}}}(k)$ is a constant [cf. Eq. (41)], we find that resulting non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ turns out to be independent of $\eta_{\mathrm{e}}$ and is given by
$b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})=\frac{k_{1}^{3}}{2\,k_{2}^{2}\,k_{3}^{2}\,(k_{2}^{3}+k_{3}^{3})}\,\left[k_{1}^{4}+k_{2}^{4}+k_{3}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+6\,k_{2}^{2}\,k_{3}^{2}\right].$ (64)

### v2

@@ IV.2 Three-point cross-correlation
We shall now turn to the computation of the three-point cross-correlation using the interaction Hamiltonian in Eq. (48). In real space, the cross-correlation between the curvature perturbation and magnetic fields is defined as
@@
where $\hat{H}_{\mathrm{int}}$ is the operator associated with the interaction Hamiltonian in Eq. (48), and $\eta_{\mathrm{i}}$ denotes the time when the initial conditions are imposed on the perturbations. Moreover, the square brackets denote the commutator between the operators denoting modes of the curvature perturbation, the magnetic field, and the Hamiltonian operator.
To describe the cross-correlation between the modes of the curvature perturbation and the magnetic field, we shall introduce the function ${\mathcal{B}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ that is defined through the relation [^3] In the definition in Eq. (52) of the three-point function, we have introduced an additional factor of $(4\,\pi)$ in comparison to the earlier works in the literature (in this context, see Refs. [90, 89, 91, 92, 103]). Note that our electromagnetic actions in Eqs. (16) and (34)) contain an additional factor of $1/(4\,\pi)$ in comparison to the action in the previous works. In order to ensure that the expression for the power spectrum of the magnetic field remains the same, we had included an overall factor of $\sqrt{4\,\pi}$ in the expression for the vector potential $A_{i}$ in Eq. (20). However, such a definition for $A_{i}$ introduces another factor of $(4\,\pi)$ when we evaluate the quantity on the right hand side of Eq. (51), which is not present if we instead follow the normalization conventions used in the aforementioned works. Therefore, in order to ensure that the quantity ${\mathcal{B}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ remains the same across the different conventions, we have introduced the factor of $(4\,\pi)$ in Eq. (52).
@@
On making use of the expression in Eq. (51), the interaction Hamiltonian in Eq. (48), the definition of the three-point function in Eq. (52) and, finally, Wick’s theorem, we find that the function ${\mathcal{B}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ can be written as follows [90, 89, 91, 92, 101, 103]:
@@
${\mathcal{G}}_{2}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$ $=2\,i\,(1+2\,n)\,(\bm{k}_{2}\cdot\bm{k}_{3})\,\int_{\eta_{\mathrm{i}}}^{\eta_{\mathrm{e}}}\frac{{\mathrm{d}}\eta}{a\,H}\,J^{2}(\eta)\,f_{k_{1}}^{\prime\ast}(\eta)\,\bar{A}_{k_{2}}^{\prime\ast}(\eta)\,\bar{A}_{k_{3}}^{\prime\ast}(\eta),$ (54b)
${\mathcal{G}}_{3}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=i\,\left[(\bm{k}_{1}\cdot\bm{k}_{2})\,k_{3}^{2}+(\bm{k}_{1}\cdot\bm{k}_{3})\,(\bm{k}_{2}\cdot\bm{k}_{3})\right]\,\int_{\eta_{\mathrm{i}}}^{\eta_{\mathrm{e}}}\frac{{\mathrm{d}}\eta}{a\,H}\,J^{2}(\eta)\,f_{k_{1}}^{\ast}(\eta)\,\bar{A}_{k_{2}}^{\prime\ast}(\eta)\,\bar{A}_{k_{3}}^{\ast}(\eta)$
$\quad+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$ (54c)
${\mathcal{G}}_{4}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{i}{k_{1}^{2}}\,\left[(\bm{k}_{1}\cdot\bm{k}_{2})\,k_{3}^{2}+(\bm{k}_{1}\cdot\bm{k}_{3})\,(\bm{k}_{2}\cdot\bm{k}_{3})\right]\,\int_{\eta_{\mathrm{i}}}^{\eta_{\mathrm{e}}}{\mathrm{d}}\eta\,J^{2}(\eta)\,\epsilon_{1}(\eta)\,f_{k_{1}}^{\prime\ast}(\eta)\,\bar{A}_{k_{2}}^{\prime\ast}(\eta)\,\bar{A}_{k_{3}}^{\ast}(\eta)$
$\quad+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$ (54d)
${\mathcal{G}}_{5}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=i\,\left[k_{2}^{2}\,k_{3}^{2}+(\bm{k}_{2}\cdot\bm{k}_{3})^{2}\right]\,\int_{\eta_{\mathrm{i}}}^{\eta_{\mathrm{e}}}{\mathrm{d}}\eta\,J^{2}(\eta)\,f_{k_{1}}^{\ast}(\eta)\,\bar{A}_{k_{2}}^{\ast}(\eta)\,\bar{A}_{k_{3}}^{\ast}(\eta),$ (54e)
${\mathcal{G}}_{6}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-i\,(1-2\,n)\,\left[k_{2}^{2}\,k_{3}^{2}+(\bm{k}_{2}\cdot\bm{k}_{3})^{2}\right]\,\int_{\eta_{\mathrm{i}}}^{\eta_{\mathrm{e}}}\frac{{\mathrm{d}}\eta}{a\,H}\,J^{2}(\eta)\,f_{k_{1}}^{\prime\ast}(\eta)\,\bar{A}_{k_{2}}^{\ast}(\eta)\,\bar{A}_{k_{3}}^{\ast}(\eta).$ (54f)
@@ IV.3 Non-Gaussianity parameter
$\hat{B}_{i\,{\bm{k}}}(\eta)$ $=\hat{B}_{i\,{\bm{k}}}^{{}_{\mathrm{G}}}(\eta)+\frac{1}{(2\,\pi)^{3/2}}\int{\mathrm{d}}^{3}\bm{p}\,\hat{{\mathcal{R}}}_{\bm{p}}^{{}_{\mathrm{G}}}(\eta)\,\hat{B}_{i\,{\bm{k}-\bm{p}}}^{{}_{\rm G}}(\eta)$
@@ V.1 Power spectra
Clearly, the amplitude of $\mathcal{P}_{{}_{\mathrm{B}}}(k)$ in ultra slow roll inflation is suppressed by a factor of $(\eta_{\mathrm{e}}/\eta_{\mathrm{i}})^{12}\sim 10^{-300}$. Such a small value of $\mathcal{P}_{{}_{\mathrm{B}}}(k)$ will not be sufficient to serve as seeds for magnetic fields with the strengths observed in the inter-galactic medium today. Despite these shortcomings of the scenario of pure ultra slow roll, we pursue the analysis with the aim of eventually utilizing the coupling function that we have constructed in a realistic model of inflation with a brief epoch of ultra slow roll. We shall now proceed with these values of the parameters to calculate and examine the amplitude and shape of the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$.
@@ V.2.1 Case of $p=6$
${\mathcal{B}}_{2}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{3\,H_{{}_{\mathrm{I}}}^{6}\,\eta_{\mathrm{i}}^{2}}{32\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{1}^{3}\,k_{2}^{3}\,k_{3}^{3}\,\epsilon_{1\mathrm{i}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{4}\,(k_{1}^{2}-k_{2}^{2}-k_{3}^{2})\,\left[k_{1}^{3}-4\,(k_{2}^{3}+k_{3}^{3})\right],$ (63b)
${\mathcal{B}}_{3}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$ $=\frac{9\,H_{{}_{\mathrm{I}}}^{6}\,\eta_{\mathrm{i}}^{2}}{128\,M_{{}_{\mathrm{Pl}}}^{2}\,\epsilon_{1\mathrm{i}}\,k_{1}^{3}\,k_{2}^{3}\,k_{3}^{5}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{4}\,\left[(k_{1}^{2}-k_{2}^{2})^{2}+2\,(k_{1}^{2}+k_{2}^{2})\,k_{3}^{2}-3\,k_{3}^{4}\right]\,(3\,k_{1}^{3}+4\,k_{2}^{3})+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$ (63c)
${\mathcal{B}}_{4}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{81\,\pi\,H_{{}_{\mathrm{I}}}^{6}}{64\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{1}^{5}\,k_{2}^{3}\,k_{3}^{5}\,\eta_{\mathrm{e}}^{3}}\,\left[(k_{1}-k_{2})^{2}+2\,(k_{1}^{2}+k_{2}^{2})\,k_{3}^{2}-3\,k_{3}^{4}\right]+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$ (63d)
${\mathcal{B}}_{5}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{9\,H_{{}_{\mathrm{I}}}^{6}}{64\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{6}\,\left[k_{1}^{4}+k_{2}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right],$ (63e)
${\mathcal{B}}_{6}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{63\,H_{{}_{\mathrm{I}}}^{6}}{64\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{6}\,\left[k_{1}^{4}+k_{2}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right].$ (63f)
Clearly, among the different contributions, it is the fifth and the sixth terms, viz. ${\cal B}_{5}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ and ${\cal B}_{6}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$, that dominate. Note that these contributions behave as $(\eta_{\mathrm{i}}/\eta_{\mathrm{e}})^{6}$, just as the scalar power spectrum $\mathcal{P}_{{}_{\mathrm{S}}}(k)$ does when $p=6$ [cf. Eq. (13)]. Since the magnetic power spectrum $\mathcal{P}_{{}_{\mathrm{B}}}(k)$ is a constant [cf. Eq. (41)], we find that resulting non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ turns out to be independent of $\eta_{\mathrm{e}}$ and is given by
$b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})=\frac{k_{1}^{3}}{2\,k_{2}^{2}\,k_{3}^{2}\,(k_{2}^{3}+k_{3}^{3})}\,\left[k_{1}^{4}+k_{2}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right].$ (64)